undeadknight26 (undeadknight26):

Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

undeadknight26 (undeadknight26):

This is what i have so far.: 2. A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 a. Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. 2 x 3 + 7 x 3 = 3 6x + 21y = 3 +-3x -4y +5 3x + 17y = 8 b. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). I used the substitution method and substituted x and y for 3. c. Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

undeadknight26 (undeadknight26):

only part C confuses me...

OpenStudy (anonymous):

lemme get this straight, these are the "orginal system" 2x + 7y = 1 -3x – 4y = 5 and the "equivalent system" is...?

OpenStudy (amistre64):

when you have a solution, plug it into the original setup ....

undeadknight26 (undeadknight26):

yes..the equivalent equation is: 3x + 17y = 8

OpenStudy (amistre64):

in order to see if x=7 is a solution to 3x=21 ... plug it in in order to show that a solution of one thing is also a solution to another .... plug em in

undeadknight26 (undeadknight26):

So how do i solve part c?

undeadknight26 (undeadknight26):

@DemolisionWolf @amistre64

OpenStudy (amistre64):

what is the solution to the equivalent eqauation?

OpenStudy (amistre64):

C is simply stating ... check to make sure youre solution works.

OpenStudy (amistre64):

A - Use an equivalent system B - Solve the equivalent system C - Make sure the solution works in the original setup. Since you already have the solutions form B, .... see if they work

undeadknight26 (undeadknight26):

i used 3 to find my equivalent equation so i have to divide them all by 3 to see if it gives me my original equations solution?

OpenStudy (amistre64):

2x + 7y = 1 -3x – 4y = 5 redo step A for me ... i cant read what youve posted. It looks like equation 1 was multiplied by 3; but what did you to to equation 2?

undeadknight26 (undeadknight26):

Nothing...was i supposed to do something to it?

OpenStudy (amistre64):

dunno, these instructions are a bit looney Multiply the first equation ... 2x + 7y = 1 6x + 21y = 3 and add it to the second equations 6x + 21y = 3 -3x – 4y = 5 -------------- 3x + 17y = 8 now we have an equivalent setup 6x + 21y = 3 3x + 17y = 8 which it says to solve by your favorite method .... which seems like an awful lot of extra work to me

OpenStudy (amistre64):

if we eliminate, multiply the bottom by -2 6x + 21y = 3 -6x -34y = -16 ---------------- -13y = -13 y = 1 replace y by 1 6x + 21(1) = 3 solve for x 6x = -18 ; x = -3 so, is x=-3, and y=1 a solution to the original setup of 2x + 7y = 1 6x + 21y = 3

undeadknight26 (undeadknight26):

ohh ok...makes sense.

OpenStudy (amistre64):

copied the wrong "original" ... 2x + 7y = 1 -3x – 4y = 5 see if they work in that setup .... also known as double checking your results

undeadknight26 (undeadknight26):

thank you for your trouble :)

OpenStudy (anonymous):

wait what is the answer to c?!?!

OpenStudy (anonymous):

@undeadknight26

OpenStudy (amistre64):

this isnt a free answering service .... what about the process did you not get?